A solid metal sphere has a radius of 3.39 cm c m and a mass of 1.877 kg k g . What is the density of the metal in g/cm3 g / c m 3 ? (The volume of sphere is V=43πr3 V = 4 3 π r 3 .)
Answer #1

Mass of sphere = 1.877kg = 1.877×1000g = 1877g

Volume of sphere = 4/3 π.r3 = 4/3 × 22/7 × (3.39cm)3

= 163.1878cm3

Density of metal = mass of sphere /volume of sphere

= 1877g/163.1878cm3

= 11.50 g/cm3

= 11.5 g/cm3. (Answer)